OpenStudy (anonymous):

You must show your work on each of the following questions. 1. Rewrite y = x^2 + 14x + 29 in general form. 2. Rewrite y = 3x^2 - 24x + 10 in general form.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

do you know what general form is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

look that up and see what you come up with.

OpenStudy (amistre64):

I agree with Hope, this is an application of a definition which can only be answered by knowing the definition.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

is for your use.

OpenStudy (anonymous):

but i dont understand it. does it mean all i have to do is plug in 0 for y and leave it?

OpenStudy (amistre64):

im pretty sure that: y = ax^2 + bx + c , is the general form

OpenStudy (anonymous):

it may be standard form^ im not positive.

OpenStudy (amistre64):

standard form is also called vertex form: y = a(x-h)^2+k

OpenStudy (anonymous):

what is h? So y = x^2 + 14x + 29 is in general form?

OpenStudy (anonymous):

yes it is in general form.

OpenStudy (amistre64):

h and k refer to the components of the vertex

OpenStudy (anonymous):

okay well what about y = 3x^2 - 24x + 10 ?

OpenStudy (anonymous):

ohh okay @amistre64

OpenStudy (anonymous):

@Hope_nicole is y = 3x^2 - 24x + 10 in general form as well?

OpenStudy (anonymous):

i dont think so, the 3 in front makes me think it isnt but im not positive. @amistre64

OpenStudy (amistre64):

its fine ... general form is: y = ax^2 + bx + c, such that a is not 0

OpenStudy (anonymous):

but for the first one a IS 0 ?

OpenStudy (amistre64):

x^2 = 1x^2 , not 0x^2

OpenStudy (anonymous):

ohh okay thank you guys so much!! @amistre64 @Hope_nicole

OpenStudy (amistre64):

you might want to reference your material that your course presents you to be sure that they are using the same definition of a general form as the rest of the sane universe :)